# Temperature profiles of accretion disks in luminous active galactic nuclei derived from ultraviolet spectroscopic variability

### v1

@@ 2.2 Sample selection
Before SF analysis, radio-loud sources ($\sim 2$% of the initial sample), classified as those with flux density ratio between 6 cm and 2500 ${\rm\AA}$ (adopted from Shen et al. 2019) higher than 10 (Kellermann et al. 1989), are eliminated because of their distinctive variability characteristics (e.g., Wagner & Witzel 1995). We also discard light curves that lack any evidence of variability, based on the variability probability criterion $P_{\rm var}<0.95$, where $P_{\rm var}$, calculated from $\chi^{2}$ distribution (McLaughlin et al. 1996; Sánchez et al. 2017), is the probability that the target is intrinsically variable. $P_{\rm var}$ is computed for each continuum window using the $\sim 7$ yr light curves and is commonly employed to confirm the intrinsic variability of the continuum flux through the $\chi^{2}$ distribution. Note that S/N2, defined as $\sqrt{\chi^{2}-{\rm DOF}}$, where DOF is the number of degrees of freedom of the light curve, is used to quantify the variability of the emission flux (Grier et al. 2019). However, we choose not to use S/N2 in this study, as the criteria for the intrinsic variability are determined through visual inspection. To compute ensemble SF in Section 3, we divide the sample by bolometric luminosity, starting from $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ = 45.4, in steps of 0.25. Objects with luminosity lower than $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 45.4$ are removed because they can be substantially contaminated by the host galaxy ($f_{\rm host}\leq 0.1$ at 3450 Å; Shen et al. 2015b), even though we only use restframe ultraviolet wavelengths. The final sample consists of 447 AGNs that have light curves in at least one continuum window among 1367, 1479, 1746, 2230, 3050, 3500, and 3800 ${\rm\AA}$. In total, 50, 85, 205, 322, 310, 249, and 197 light curves are available, respectively, for the continuum windows at 1367, 1479, 1746, 2230, 3050, 3500, and 3800 ${\rm\AA}$.
@@ 3 Ensemble structure function
Following the definition from Press et al. (1992), the SF of an individual object is calculated as
@@
where $N_{\Delta t,{\rm pair}}$ is the number of observed pairs with the time lag ($\Delta t$), $m$ is the observed magnitude, and $\sigma_{\rm final}$ is the final uncertainty of magnitude in the single-epoch spectrum, as described in Section 2.1. We adopt this SF definition because incomplete noise terms often lead to a flatter SF than the real SF (Kozlowski 2016). Because the SF can be highly sensitive to outliers, we apply $3\,\sigma$ clipping to the median with half of the difference between 16th and 84th percentile of magnitude measurements in each light curve.
To examine the average SF as a function of luminosity and wavelength, we choose to use an ensemble SF instead of the SF from individual objects. The ensemble SF is calculated by averaging SF2s of sources in each luminosity and wavelength bin, considering that the SFs of individual sources are often associated with systematically large uncertainties and have negative values (e.g., Son et al. 2023; Kim et al. 2024). To avoid potential bias due to the small sample size, the ensemble SF is computed only if there are at least 10 objects in a group. Table 1 gives the number of objects used to estimate the ensemble SF and AGN properties for each group. The derived ensemble SFs in different luminosity and wavelength bins are shown in Figure 3.
To evaluate potential biases in estimating the ensemble SFs caused by the seasonal gaps and the limited baseline in the light curves, we perform the simulations using regularly sampled light curves with a cadence of one day, assuming the damped random walk model. Subsequently, we generate mock light curves, incorporating the actual cadence and redshift distributions of our sample. This experiment demonstrates that the ensemble SFs for $\Delta t\lessapprox 200$ days can be reproduced accurately using the simulated light curves (see also Emmanoulopoulos et al. 2010; Kozlowski 2017).
@@ 4 Results
To understand the physical origin of the variability, we examine how the characteristic timescale depends on the physical properties of AGNs. Previous studies demonstrated that $\tau$ might correspond to the thermal timescale (Kelly et al. 2009, 2013; Caplar et al. 2017; Tang et al. 2023; Arévalo et al. 2024). In the standard disk model (Shakura & Sunyaev 1973; Frank et al. 2002), the thermal timescale correlates with the orbital timescale ($t_{\rm orb}$):
@@
where $R_{\lambda}$ is the characteristic disk radius responsible for continuum emission at the specific restframe wavelength, and $\alpha$ is the viscosity parameter, assumed constant in the standard disk model. Since the disk radiates blackbody emission, its effective temperature ($T$) at a given disk radius is given by
@@
From Equations (2) and (4) with an assumption of constant $\alpha$ and $L\propto\dot{M}$, the thermal timescale is correlated with the luminosity and restframe wavelength ($\lambda^{2}_{\rm RF}$) as
@@
To test whether this holds true for our dataset, we examine the dependence of $\tau$ on the AGN luminosity and wavelength in the form
@@
The length of the light curves is insufficiently long to constrain $\tau$ directly. Instead, we adopt an indirect method to investigate the dependence of $\tau$ on the AGN luminosity and wavelength, assuming that all the SFs follow a single universal function if $\Delta t$ is converted to units of $\tau$ (e.g., Tang et al. 2023; Arévalo et al. 2024; Petrecca et al. 2024) as
@@
where $\gamma$ is the power-law slope of SF and $\delta$ is the SF at $\Delta t=\tau$. Since $\tau$ is derived in arbitrary units, the value of $\delta$ also has no physical meaning in this study. Fitting the ensemble SFs of each luminosity group at different wavelengths, we can investigate the dependence of $\tau$ on luminosity and wavelength. To achieve this, we first compute the ensemble SFs within luminosity bins for each continuum window and then fit them using the Python package lmfit to determine $a$ and $b$.
We derive $a=0.50\pm 0.03$ and $b=1.42\pm 0.09$ (Figure 3b). Whereas the best-fit value of $a$ is in excellent agreement with the predicted value of 0.5 from the standard disk model, the value of $b$ is substantially smaller than the predicted value of 2, by more than $6\,\sigma$ level. This reveals that the temperature (color) radial profile of the accretion disk deviates from the standard disk.
Figure 3: Ensemble SF as a function of (a) $\Delta t$ and (b) $\Delta t/\tau$ at different wavelengths for each group color-coded by the average bolometric luminosity. Different symbols indicate different wavelengths used for continuum extraction. Panel (b) assumes that all the SFs follow a single universal function if $\Delta t$ is converted to units of $\tau\propto L_{\rm bol}^{a}\lambda_{\rm RF}^{b}$. Note that values of $\Delta t/\tau$ on the x-axis are only meaningful in relative terms. The solid black line represents the best-fit universal SF.
@@ 5 Discussion
5.1 Possible origins of distinctive color profile
@@
Here, we only focus on the impact of slim disks, as they account for nearly one-quarter of the final sample; highly sub-Eddington systems in the RIAF state account for less than 1% of our sources. Considering that a slim disk has a distinctive geometry, temperature profile, and kinematics compared to the standard disk, its characteristic timescale can differ from that of a standard disk (Beloborodov 1998; Wang & Zhou 1999; Watarai et al. 2000). However, the state change mostly occurs inside the photon-trapping radius of the slim disk, where the accretion disk becomes vertically extended due to the radiation pressure (Abramowicz et al. 1988; Wang & Zhou 1999; Sadowski 2011). To test if this phenomenon is directly related to our finding, we first compute the photon-trapping radius and compare it with the effective emitting radius corresponding to the continuum wavelengths used in this study. The self-similar solution of Wang & Zhou (1999) for a slim disk expresses the photon-trapping radius as
@@
where $\dot{m}=\dot{M}/\dot{M}_{\rm Edd}$ is dimensionless accretion rate and $R_{g}=2GM/c^{2}$ is the Schwarzschild radius. In the same model, the effective temperature at a given radius $R$ is given by
$T_{\rm eff}=2.1\times 10^{7}\,(M_{\rm BH}/M_{\odot})^{-1/4}(R/R_{g})^{-1/2}.$ (9)
@@
We note that $\dot{m}$ cannot be directly measured for our sample because the BH spin and the viscous parameter are unknown. Therefore, we conservatively assume that the BH spin is 0 and the viscous parameter is 0.01, whereupon $\dot{m}$ is highest and hence the photon-trapping radius is largest at a given Eddington ratio. With these assumptions, we estimate $\dot{m}$ visually from Figure 4.11 of Sadowski (2011); the uncertainty of this visual estimate is $\sim 3\%$ ($0.01$ dex). With the estimate of $\dot{m}$ in hand, Equation (10) can then be used to deduce that the photon-trapping wavelength $\lambda_{\rm tr}$ for the majority of the sample ranges from $\sim 300$ to $\sim 1350$, which is significantly shorter than the continuum wavelengths ($1367-3800{\rm\AA}$) used in this study [^5] The only exception is a single object for which $\lambda_{\rm tr}=1482\,{\rm\AA}$ is larger than 1367 and 1479 ${\rm\AA}$. Omitting this target hardly changes the fit of Equation (6).. The continuum fluxes analyzed here mostly originate from emitting regions outside the photon-trapping radius for highly accreting AGNs. It is also noteworthy that, from both theoretical (Watarai et al. 2000) and observational considerations (Li et al. 2024), the effective temperature profile of a slim disk outside of its photon-trapping radius is not markedly different from that of a standard thin disk. We conclude that the discrepancy of the $\tau-\lambda_{\rm RF}^{b}$ between this study ($b=1.43\pm 0.09$) and the prediction from the standard disk ($b=2$) is unlikely to originate from the transition to the slim disk.
@@
To test the impact of RIAFs, we additionally reject the four AGNs with RIAFs. The fit to Equation (6) remains almost the same ($a=0.49\pm 0.03$, $b=1.43\pm 0.09$). As a final sanity check, we apply the fit to the ensemble SFs for objects with $0.01<\lambda_{\rm Edd}\leq 0.3$ that strictly speaking conform to the expectations of a standard disk, and still the results remain the same ($a=0.46\pm 0.07$, $b=1.32\pm 0.15$; Table 2). Neither slim disks nor RIAFs have a significant impact on our results.
@@
Another factor that might contribute to the departure from the expected color profile is contamination from the Balmer pseudocontinuum originating from the BLR, which contributes $\sim 10\%$ on average to the flux density at 3000 ${\rm\AA}$ (Kovačević-Dojčinović et al. 2017). This pseudo-continuum has also been regarded as a culprit for the disk size problem identified from continuum RM studies (e.g., Fausnaugh et al. 2016; McHardy et al. 2018). To evaluate the possible effect of the Balmer continuum on our results, we refit Equation (6) by excluding wavelengths longer than 3000 ${\rm\AA}$, the spectral range most severely contaminated by the Balmer continuum (e.g., Cackett et al. 2018). The fitting results hardly change ($a=0.50\pm 0.04$, $b=1.45\pm 0.18$) within a $1\,\sigma$ level. Therefore, the Balmer continuum does not seem to be the origin of the color profile discrepancy.
@@ 5.2 Comparison with observational studies
Various studies have investigated the color profiles in luminous AGNs based on light curves from broadband photometric data. For example, using the broadband data from the first year of the SDSS-RM project, Homayouni et al. (2019) found that the color profile ($R\propto\lambda_{\rm RF}^{p}$) is well fit with $p=1.28_{-0.39}^{+0.41}$, equivalent to $b=3p/2=1.92_{-0.585}^{+0.615}$. Their estimation is consistent with both the standard disk prediction and our best fit, considering their considerably larger uncertainties. Tang et al. (2023), employing the same method as ours, showed that the characteristic timescale is proportional to $L^{0.539\pm 0.004}\lambda^{2.418\pm 0.023}$; their values of $a$ and $b$ are formally steeper than those predicted by the standard disk. It is unclear why their estimate of $b$ is significantly larger than that of our study. We speculate that the discrepancy may arise from a potential bias introduced by the broadband photometry data. In addition, Tang et al. (2023) used a longer $\Delta t$ (up to 350 days) compared to our study (up to 179 days). We impose this $\Delta t$ limit because of the restricted baseline of our dataset (e.g., Emmanoulopoulos et al. 2010), which is more pronounced at shorter wavelengths (1367 and 1479 Å) because of cosmological time dilation. However, two studies based on photometric data offer counter-examples. Analyzing the characteristic timescales of thousands of individual quasars using $\sim 10$ years of broadband light curves from the SDSS Stripe 82 region (Jiang et al. 2014), MacLeod et al. (2010) found $\tau\propto\lambda_{\rm RF}^{0.17\pm 0.02}$. Similarly, Stone et al. (2022) (see erratum in Stone et al. 2023) studied $\sim 20$ years of broadband light curves for 190 SDSS Stripe 82 quasars and found $\tau\propto\lambda_{\rm RF}^{0.35\pm 0.14}$ or $\lambda_{\rm RF}^{0.43\pm 0.01}$ depending on the sample selection. These works produced significantly shallower slopes than our study. The situation for continuum RM based on spectroscopic data is no less confusing. While the work of Fausnaugh et al. (2016) and Starkey et al. (2017) yielded $p=0.99\pm 0.14$ or $b=1.49\pm 0.21$, which agrees well with our best-fit results, most are at variance, with a preponderance of studies supporting $p\approx 4/3$ (Cackett et al. 2018; Collier et al. 1998; Homayouni et al. 2019; Guo et al. 2022).
@@
Microlensing analysis, too, has produced a wide diversity of power-law indices, ranging from $p\approx 4/3$ (Eigenbrod et al. 2008; Poindexter et al. 2008; Rivera et al. 2023) to $p<4/3$ (Floyd et al. 2009; Blackburne et al. 2011; Jiménez-Vicente et al. 2014; Muñoz et al. 2016) and $p>4/3$ (Bate et al. 2018; Cornachione & Morgan 2020). It is noteworthy that microlensing studies based on spectroscopic data and/or narrowband photometric data have a tendency to favor smaller values of the power-law index ($p\approx 0.7-1.3$ or $b\approx 1.05-1.95$; Eigenbrod et al. 2008; Mediavilla et al. 2011; Mosquera et al. 2011; Motta et al. 2012; Jiménez-Vicente et al. 2014; Muñoz et al. 2016) compared to the prediction from the standard disk model, which qualitatively aligns with our study.
@@
5.3 Application to accretion disk model
Assuming that AGNs obey a universal SF, we find that the characteristic timescale ($\tau$) is proportional to $L^{0.5}$ and $\lambda_{\rm RF}^{1.42}$. While the luminosity dependence is consistent with the prediction from the standard disk theory, the wavelength scaling reveals that the radial temperature profile is considerably steeper than that of a standard disk. After careful consideration of factors that might affect the analysis, we conclude that this discrepancy in the temperature profile is likely to be genuine.
There is presently no obvious resolution to this dilemma. From a theoretical point of view, a steep temperature (and hence shallow color) profile can be partly explained by the magneto-rotational instability model attributed to additional dissipation originating from torque at the inner edge of the accretion disk. For example, the magneto-rotational instability model of Agol & Krolik (2000) predicts a power-law index of $p=8/7=1.14$ ($b=12/7=1.71$), which is marginally smaller than $p=4/3$ ($b=2$), although not as small as our observed value of $b=1.42$. On the other hand, the plethora of scenarios that have been proposed to solve the disk size problem either do not focus on the temperature profile or they predict temperature profiles contradicting our results (e.g., Lawrence 2012; Gardner & Done 2017; Kammoun et al. 2019; Li et al. 2019; Sun et al. 2019, 2020; Starkey et al. 2023). For instance, Lawrence (2012) argued that reprocessed continuum emission, presumably due to cold, dense clouds located at $\sim 30\,R_{S}$, may be crucial to explain the disk size problem. Contrary to our results, however, a steeper temperature profile is expected from this model. In a similar manner, the disk wind model favors not only a steeper temperature profile but also a larger power-law index for the AGN luminosity (i.e., $a>0.5$; Li et al. 2019; Sun et al. 2019). The temperature profile of the rimmed accretion disk model of Starkey et al. (2023) does not depart substantially (i.e., $T\propto R^{-3/4}$) from that of the standard disk except for the outer part of the disk.

### v2

@@ 2.2 Sample selection
Before SF analysis, radio-loud sources ($\sim 2$% of the initial sample), classified as those with flux density ratios between 6 cm and 2500 ${\rm\AA}$ (adopted from Shen et al. 2019) higher than 10 (Kellermann et al. 1989), were eliminated because of their distinctive variability characteristics (e.g., Wagner & Witzel 1995). We also discarded light curves that lack evidence of variability based on the variability probability criterion $P_{\rm var}<0.95$, where $P_{\rm var}$, calculated from the $\chi^{2}$ distribution (McLaughlin et al. 1996; Sánchez et al. 2017), is the probability that the target is intrinsically variable. We computed $P_{\rm var}$ for each continuum window using the $\sim 7$ yr light curves; $P_{\rm var}$ is commonly employed to confirm the intrinsic variability of the continuum flux through the $\chi^{2}$ distribution. We note that S/N2, defined as $\sqrt{\chi^{2}-{\rm DOF}}$, where DOF is the number of degrees of freedom of the light curve, is typically used to quantify the variability of the emission flux (Grier et al. 2019). However, we chose not to use S/N2 in this study, as the criteria for the intrinsic variability were determined through visual inspection. To compute ensemble SFs (see Sect. 3), we divided the sample based on the bolometric luminosity,starting from $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ = 45.4 in steps of 0.25. Objects with luminosities lower than $\log\ (L_{\rm bol}/{\rm erg\ s^{-1}})$ $\leq 45.4$ were removed because they can be substantially contaminated by the host galaxy ($f_{\rm host}\leq 0.1$ at 3450 Å; Shen et al. 2015b), even though we only use rest-frame ultraviolet wavelengths. The final sample consists of 447 AGNs that have light curves in at least one continuum window among 1367, 1479, 1746, 2230, 3050, 3500, and 3800 ${\rm\AA}$. In total, 50, 85, 205, 322, 310, 249, and 197 light curves are available, respectively, for the continuum windows at 1367, 1479, 1746, 2230, 3050, 3500, and 3800 ${\rm\AA}$.
@@ 3 Ensemble structure function
Following the definition from Press et al. (1992), the SF of an individual object was calculated as
@@
where $N_{\Delta t,{\rm pair}}$ is the number of observed pairs with the time lag ($\Delta t$), $m$ is the observed magnitude, and $\sigma_{\rm final}$ is the final uncertainty of magnitude in the single-epoch spectrum, as described in Sect. 2.1. We adopted this SF definition because incomplete noise terms often lead to a flatter SF than the real SF (Kozlowski 2016). Because the SF can be highly sensitive to outliers, we applied $3\,\sigma$ clipping to the median with half of the difference between the 16th and 84th percentile of magnitude measurements in each light curve.
To examine the average SF as a function of luminosity and wavelength, we used an ensemble SF instead of the SF from individual objects. The ensemble SF was calculated by averaging SF2s of sources in each luminosity and wavelength bin, considering that the SFs of individual sources are often associated with systematically large uncertainties and have negative values (e.g., Son et al. 2023; Kim et al. 2024). To avoid potential bias due to the small sample size, the ensemble SF was computed only if there were at least ten objects in a group. The number of objects used to estimate the ensemble SF and AGN properties for each group is provided in Appendix A. The derived ensemble SFs in different luminosity and wavelength bins are shown in Fig. 3.
To evaluate potential biases in estimating the ensemble SFs caused by the seasonal gaps and the limited baseline in the light curves, we performed the simulations using regularly sampled light curves with a cadence of one day, assuming the damped random walk model. Subsequently, we generated mock light curves, incorporating the actual cadence and redshift distributions of our sample. This experiment demonstrates that the ensemble SFs for $\Delta t\lessapprox 200$ days can be reproduced accurately using the simulated light curves (see also Emmanoulopoulos et al. 2010; Kozlowski 2017).
@@ 4 Results
To understand the physical origin of the variability, we examined how the characteristic timescale depends on the physical properties of AGNs. Previous studies demonstrated that $\tau$ might correspond to the thermal timescale (Kelly et al. 2009, 2013; Caplar et al. 2017; Tang et al. 2023; Arévalo et al. 2024). In the standard disk model (Shakura & Sunyaev 1973; Frank et al. 2002), the thermal timescale correlates with the orbital timescale ($t_{\rm orb}$):
@@
where $R_{\lambda}$ is the characteristic disk radius responsible for continuum emission at the specific rest-frame wavelength, and $\alpha$ is the viscosity parameter, assumed constant in the standard disk model. Since the disk radiates blackbody emission, its effective temperature ($T$) at a given disk radius is given by
@@
From Eqs. (2) and (4) with an assumption of constant $\alpha$ and $L\propto\dot{M}$, the thermal timescale is correlated with the luminosity and rest-frame wavelength ($\lambda^{2}_{\rm RF}$) as
@@
To test whether this holds true for our dataset, we examined the dependence of $\tau$ on the AGN luminosity and wavelength in the form
@@
The light curves are insufficiently long to constrain $\tau$ directly. Instead, we adopted an indirect method to investigate the dependence of $\tau$ on the AGN luminosity and wavelength, assuming that all the SFs follow a single universal function if $\Delta t$ is converted to units of $\tau$ (e.g., Tang et al. 2023; Arévalo et al. 2024; Petrecca et al. 2024) as
@@
where $\gamma$ is the power-law slope of the SF and $\delta$ is the SF at $\Delta t=\tau$. Since $\tau$ is derived in arbitrary units, the value of $\delta$ has no physical meaning in this study. Fitting the ensemble SFs of each luminosity group at different wavelengths, we can investigate the dependence of $\tau$ on luminosity and wavelength. To achieve this, we first computed the ensemble SFs within luminosity bins for each continuum window and then fit them using the Python package lmfit to determine $a$ and $b$.
We derive $a=0.50\pm 0.03$ and $b=1.42\pm 0.09$ (Fig. 3b). While the best-fit value of $a$ is in excellent agreement with the predicted value of 0.5 from the standard disk model, the value of $b$ is substantially smaller than the predicted value of 2, by more than $6\,\sigma$. This reveals that the temperature (color) radial profile of the accretion disk deviates from the standard disk model.
Figure 3: Ensemble SF as a function of $\Delta t$ (a) and $\Delta t/\tau$ (b) at different wavelengths for each group, color-coded by the average bolometric luminosity. Different symbols indicate the different wavelengths used for continuum extraction. Panel (b) assumes that the SFs all follow a single universal function if $\Delta t$ is converted to units of $\tau\propto L_{\rm bol}^{a}\lambda_{\rm RF}^{b}$. Note that the values of $\Delta t/\tau$ on the x-axis are only meaningful in relative terms. The solid black line represents the best-fit universal SF.
@@ 5 Discussion
5.1 Possible origins of the distinctive color profile
@@
Here, we only focused on the impact of slim disks, as they account for nearly one-quarter of the final sample; highly sub-Eddington systems in the RIAF state account for less than 1% of our sources. Considering that a slim disk has a distinctive geometry, temperature profile, and kinematics compared to the standard disk, its characteristic timescale can differ from that of a standard disk (Beloborodov 1998; Wang & Zhou 1999; Watarai et al. 2000). However, the state change mostly occurs inside the photon-trapping radius of the slim disk, where the accretion disk becomes vertically extended due to the radiation pressure (Abramowicz et al. 1988; Wang & Zhou 1999; Sadowski 2011). To test if this phenomenon is directly related to our finding, we first computed the photon-trapping radius and compared it with the effective emitting radius corresponding to the continuum wavelengths used in this study. The self-similar solution of Wang & Zhou (1999) for a slim disk expresses the photon-trapping radius as
@@
where $\dot{m}=\dot{M}/\dot{M}_{\rm Edd}$ is the dimensionless accretion rate and $R_{g}=2GM/c^{2}$ is the Schwarzschild radius. In the same model, the effective temperature at a given radius ($R$) is given by
$T=2.1\times 10^{7}\,(M_{\rm BH}/M_{\odot})^{-1/4}(R/R_{g})^{-1/2}.$ (9)
@@
We note that $\dot{m}$ cannot be directly measured for our sample because the BH spin and the viscous parameter are unknown. Therefore, we conservatively assumed that the BH spin is 0 and the viscous parameter is 0.01, whereupon $\dot{m}$ is highest and hence the photon-trapping radius is largest at a given Eddington ratio. With these assumptions, we estimated $\dot{m}$ visually from Fig. 4.11 of Sadowski (2011); the uncertainty of this visual estimate is $\sim 3\%$ ($0.01$ dex). With the estimate of $\dot{m}$ in hand, Eq. (10) can then be used to deduce that the photon-trapping wavelength ($\lambda_{\rm tr}$) for the majority of the sample ranges from $\sim 300$ to $\sim 1350$, which is significantly shorter than the continuum wavelengths ($1367-3800{\rm\AA}$) used in this study [^5] The only exception is a single object for which $\lambda_{\rm tr}=1482\,{\rm\AA}$ is larger than 1367 and 1479 ${\rm\AA}$, and omitting this target hardly changes the fit of Eq. (6).. The continuum fluxes analyzed here mostly originate from emitting regions outside the photon-trapping radius for highly accreting AGNs. It is also noteworthy that, based on both theoretical (Watarai et al. 2000) and observational considerations (Li et al. 2024), the effective temperature profile of a slim disk outside of its photon-trapping radius is not markedly different from that of a standard thin disk. We conclude that the discrepancy of the $\tau-\lambda_{\rm RF}^{b}$ between this study ($b=1.43\pm 0.09$) and the prediction from the standard disk model ($b=2$) is unlikely to originate from the transition to the slim disk.
@@
To test the impact of RIAFs, we additionally rejected the four AGNs with RIAFs. The fit to Eq. (6) remains almost the same ($a=0.49\pm 0.03$, $b=1.43\pm 0.09$). As a final sanity check, we applied the fit to the ensemble SFs for objects with $0.01<\lambda_{\rm Edd}\leq 0.3$ that strictly speaking conform to the expectations of a standard disk, and still the results remain the same ($a=0.46\pm 0.07$, $b=1.32\pm 0.15$; Table 1). Neither slim disks nor RIAFs have a significant impact on our results.
@@
Another factor that might contribute to the departure from the expected color profile is contamination from the Balmer pseudo-continuum originating from the BLR, which contributes $\sim 10\%$ on average to the flux density at 3000 ${\rm\AA}$ (Kovačević-Dojčinović et al. 2017). This pseudo-continuum has also been regarded as a culprit for the disk size problem identified from continuum RM studies (e.g., Fausnaugh et al. 2016; McHardy et al. 2018). To evaluate the possible effect of the Balmer continuum on our results, we refit Eq. (6) by excluding wavelengths longer than 3000 ${\rm\AA}$, the spectral range most severely contaminated by the Balmer continuum (e.g., Cackett et al. 2018). The fitting results hardly change ($a=0.50\pm 0.04$, $b=1.45\pm 0.18$) within a $1\,\sigma$ level. Therefore, the Balmer continuum does not seem to be the origin of the color profile discrepancy either.
Table 1: Best-fit parameters for $\tau\propto L^{a}\,\lambda^{b}_{\rm RF}$.
Subsample $a$ $b$
(1) (2) (3)
All $0.50\pm 0.03$ $1.42\pm 0.09$
$0.01<\lambda_{\rm Edd}\leq 0.3$ $0.46\pm 0.07$ $1.32\pm 0.15$
All with $\lambda_{\rm RF}<3000\,{\rm\AA}$ $0.50\pm 0.04$ $1.45\pm 0.18$
@@ 5.2 Comparison with observational studies
Various studies have investigated the color profiles in luminous AGNs based on light curves from broadband photometric data. For example, using the broadband data from the first year of the SDSS-RM project, Homayouni et al. (2019) found that the color profile ($R\propto\lambda_{\rm RF}^{p}$) is well fit with $p=1.28_{-0.39}^{+0.41}$, equivalent to $b=3p/2=1.92_{-0.585}^{+0.615}$. Their estimation is consistent with both the standard disk prediction and our best fit, considering their considerably larger uncertainties. Tang et al. (2023), employing the same method as ours, showed that the characteristic timescale is proportional to $L^{0.539\pm 0.004}\lambda^{2.418\pm 0.023}$; their values of $a$ and $b$ are formally steeper than those predicted by the standard disk. It is unclear why their estimate of $b$ is significantly larger than that of our study. We speculate that the discrepancy may arise from a potential bias introduced by the broadband photometry data. In addition, Tang et al. (2023) used a longer $\Delta t$ (up to 350 days) than we did (up to 179 days). We imposed this $\Delta t$ limit because of the restricted baseline of our dataset (e.g., Emmanoulopoulos et al. 2010), which is more pronounced at shorter wavelengths (1367 and 1479 Å) because of cosmological time dilation. However, two studies based on photometric data offer counter-examples. Analyzing the characteristic timescales of thousands of individual quasars using $\sim 10$ years of broadband light curves from the SDSS Stripe 82 region (Jiang et al. 2014), MacLeod et al. (2010) found $\tau\propto\lambda_{\rm RF}^{0.17\pm 0.02}$. Similarly, Stone et al. (2022, see the erratum in ) studied $\sim 20$ years of broadband light curves for 190 SDSS Stripe 82 quasars and found $\tau\propto\lambda_{\rm RF}^{0.35\pm 0.14}$ or $\lambda_{\rm RF}^{0.43\pm 0.01}$ depending on the sample selection. These works produced significantly shallower slopes than our study. The situation for continuum RM based on spectroscopic data is no less confusing. While the work of Fausnaugh et al. (2016) and Starkey et al. (2017) yielded $p=0.99\pm 0.14$ or $b=1.49\pm 0.21$, which does agree well with our best-fit results, most are at variance, with a preponderance of studies supporting $p\approx 4/3$ (Cackett et al. 2018; Collier et al. 1998; Homayouni et al. 2019; Guo et al. 2022).
@@
Microlensing analysis, too, has produced a wide diversity of power-law indices, ranging from $p\approx 4/3$ (Eigenbrod et al. 2008; Poindexter et al. 2008; Rivera et al. 2023) to $p<4/3$ (Floyd et al. 2009; Blackburne et al. 2011; Jiménez-Vicente et al. 2014; Muñoz et al. 2016) and $p>4/3$ (Bate et al. 2018; Cornachione & Morgan 2020). It is noteworthy that microlensing studies based on spectroscopic data and/or narrowband photometric data have a tendency to favor smaller power-law index values ($p\approx 0.7-1.3$ or $b\approx 1.05-1.95$; Eigenbrod et al. 2008; Mediavilla et al. 2011; Mosquera et al. 2011; Motta et al. 2012; Jiménez-Vicente et al. 2014; Muñoz et al. 2016) compared to the prediction from the standard disk model; this qualitatively aligns with our study.
@@
5.3 Application to the accretion disk model
Assuming that AGNs obey a universal SF, we find that the characteristic timescale ($\tau$) is proportional to $L^{0.5}$ and $\lambda_{\rm RF}^{1.42}$. While the luminosity dependence is consistent with the prediction from standard disk theory, the wavelength scaling reveals that the radial temperature profile is considerably steeper than that of a standard disk. After careful consideration of factors that might affect the analysis, we conclude that this discrepancy in the temperature profile is likely genuine.
There is presently no obvious resolution to this dilemma. From a theoretical point of view, a steep temperature (and hence a shallow color) profile can be partly explained by the magnetorotational instability model attributed to additional dissipation originating from torque at the inner edge of the accretion disk. For example, the magnetorotational instability model of Agol & Krolik (2000) predicts a power-law index of $p=8/7=1.14$ ($b=12/7=1.71$), which is marginally smaller than $p=4/3$ ($b=2$), although not as small as our observed value of $b=1.42$. The steep temperature profile can also be produced by the continuum emission at a specific wavelength, which is contaminated by emission from different radii (Zhou et al. 2024). Considering this contamination, Zhou et al. (2024) demonstrated that $a$ and $b$ can be 0.65 and 1.19, respectively, from the simulated mock light curves based on the corona-heated accretion-disk reprocessing (CHAR) model (Sun et al. 2020). While their $a$ is marginally higher than our result, their predicted value of $b$ broadly aligns with our observation within $3\sigma$.
On the other hand, the plethora of scenarios that have been proposed to solve the disk size problem either do not focus on the temperature profile or predict temperature profiles that contradict our results (e.g., Lawrence 2012; Gardner & Done 2017; Kammoun et al. 2019; Li et al. 2019; Sun et al. 2019, 2020; Starkey et al. 2023). For instance, Lawrence (2012) argued that reprocessed continuum emission, presumably due to cold, dense clouds located at $\sim 30\,R_{S}$, may be crucial to explain the disk size problem. Contrary to our results, however, a steeper temperature profile is expected from this model. In a similar manner, the disk wind model favors not only a steeper temperature profile but also a higher power-law index for the AGN luminosity (i.e., $a>0.5$; Li et al. 2019; Sun et al. 2019). The temperature profile of the rimmed accretion disk model of Starkey et al. (2023) does not depart substantially (i.e., $T\propto R^{-3/4}$) from that of the standard disk, except for the outer part of the disk.